What is the dose in grams for high dose methotrexate 5 g/m2 is for a child who weighs 66 lb and is 47.2 inches in height?
Answer #1

According to Du Bois formula for calculating BSA (Body Surface Area) = 0.007184 * W0.425 * H0.725

where W is the weight in Kg and H the height in cm.

Given, Weight of the child = 66 lb

To convert lb to Kg, we divide the lb by 2.205: 66/2.205 = 29.9 Kg

Height = 47.2 inches.

1 inch = 2.5 cm. Therefore , 47.2*2.5 = 118 cm.

BSA of the child = 0.007184* 29.90.425 * 1180.725

= 0.007184* 4.27* 31.8

= 0.975 m2

Now, for a dosage of 5 g/ m2, the amount of methotrexate to be administered to the child would be : 0.975*5 = 4.875 g.